OpenStudy (anonymous):

Hey guys please help me out.. "The compound interest,calculated yearly on a certain sum of money for the second year is Rs. 880 and for the third year it is Rs.968. Calculate the rate of interest and the sum of money.. Administrators...please help..

OpenStudy (anonymous):

Can you please show me the working which mifht help me to understand... Then i can clear my doubts and have a clear concept @campbell_st!!!! Pl help..

OpenStudy (anonymous):

Um.. Actually i dont think the time period is 1year. Because i mentioned in the question that 880 is theC.I. For the 2nd year so i guess the time is 2years? @campbell_st

OpenStudy (anonymous):

Oh well i'll try and let you know in a bit... Thank you so much..

OpenStudy (anonymous):

Well @campbell_st thank u sooo.... Much... The answer is 11% right? Just confirm...

OpenStudy (anonymous):

Yea... I'm sorry dat ws a typing error its 10% but they've asked us to find the sum of money as well so how would i do that? @campbell_st

OpenStudy (anonymous):

Well is the answer 8000?

OpenStudy (anonymous):

@campbell_st?

OpenStudy (anonymous):

But according to my book it iis Rupees 8000... And i didnt understand. Why the C.I for the second year is taken as the principal.. Plz explain ot once again...

OpenStudy (anonymous):

@campbell_st??!

OpenStudy (anonymous):

Yea... Although the rate % is correct..

OpenStudy (anonymous):

Okay so theres an equation like P=968000/331? Right?

OpenStudy (anonymous):

@campbell_st??

OpenStudy (anonymous):

Oh well i'm sorry for disturbing you late at night.. Its morning here though... Anyways thank you so much for helping me till this far....@campbell_st

OpenStudy (anonymous):

Can any one else please help? @satellite73

OpenStudy (kropot72):

Let the principal = P Let the annual rate of interest as a decimal = r The amount at the end of the first year including interest = P(1 + r) The amount at the end of the second year = P(1 + r) + {(P(1 + r) * r)} The amount at the end of the third year = P(1 + r) + {(P(1 + r) * r)} + P(1 + r) + [{(P(1 + r) * r)} *r] We can now form two equations using the given amounts of interest: {(P(1 + r) * r)} = 880 ......................................(1) {(P(1 + r) * r)} + {(P * r^2)(1 + r)} = 968 .........(2) Substituting the value of {(P(1 + r) * r)} obtained in equation (1) into equation (2) gives {(P * r^2)(1 + r)} = 88 which gives \[P=\frac{88}{r ^{2}(1+r)}\ ....................(3)\] Substituting the value for P obtained in equation (3) into equation (1) we get \[\frac{88r}{r ^{2}}=880\] \[\frac{88}{r}=880\] r = 0.1 Substituting this value for r into equation (1) we find that P = 8000

OpenStudy (anonymous):

thank u @kropot72 !!! thanks a zillion..

OpenStudy (kropot72):

You're welcome :)

OpenStudy (anonymous):

@kropot72 can u say y v need to add d interest in d second year?

OpenStudy (anonymous):

@kropot72 ??

OpenStudy (kropot72):

The interest is added at the end of the second year: The amount at the end of the second year = P(1 + r) + {(P(1 + r) * r)} the interest being {(P(1 + r) * r)}

OpenStudy (anonymous):

still dont get it!

OpenStudy (anonymous):

@kropot72 ???

OpenStudy (kropot72):

The amount at the end of the first year including interest = P(1 + r) Do you understand the above?

OpenStudy (anonymous):

yes.. proceed..

OpenStudy (kropot72):

The interest for the second year is calculated on the amount at the end of the first year. The amount at the end of the first year including interest = P(1 + r) The interest on P(1 + r) for a year is {(P(1 + r) * r)} being the amount at the end of the first year multiplied by the interest rate.

OpenStudy (anonymous):

oh... thanks....now i got it.. thank u and gudni8! @kropot72

OpenStudy (kropot72):

You're welcome :)